Text Problem 25-5 each leel of cutput, begiring with the first urit (Eavakes nounded o the nearet penny) Book Costs 2.00 7.75 5.25 7 50 Ravcnuc (S Protit RevnueICost (S 8.00 750 6.50 6 25 5.80 19.75 24.50
Answer #1
Q P TC TR TP MR MC
0 8 2 0 -2
1 7.75 5.25 7.75 2.5 7.75 3.25
2 7.5 7.5 15 7.5 7.25 2.25
3 7.25 9.6 21.75 12.15 6.75 2.1
4 7 12.1 28 15.9 6.25 2.5
5 6.75 15 33.75 18.75 5.75 2.9
6 6.5 19.75 39 19.25 5.25 4.75
7 6.25 24.5 43.75 19.25 4.75 4.75

TR = P x Q

TP = TR - TC

MR (nth unit) = TR (n units) - TR ((n-1) units)

MC (nth unit) = TC (n units) - TC ((n-1) units)